# Accuracy of Charge Densities in Electronic Structure Calculations

### v3

@@ Abstract
Accurate charge densities are essential for reliable electronic structure calculations because they significantly impact predictions of various chemical properties and in particular, according to the Hellmann-Feynman theorem, atomic forces. This study examines the accuracy of charge densities obtained from different DFT exchange-correlation functionals in comparison with coupled cluster calculations with single and double excitations. We find that modern DFT functionals can provide highly accurate charge densities, particularly in case of meta-GGA and hybrid functionals. In connection with Gaussian basis sets, it is necessary to use the largest basis sets available to obtain densitites that are nearly basis set error free. These findings highlight the importance of selecting appropriate computational methods for generating high-precision charge densities, which are for instance needed to generate reference data for training modern machine learned potentials.
@@ I Introduction
Obtaining an accurate charge density is essential for predicting chemical properties, making it a cornerstone in quantum chemistry and materials science. The Hellmann-Feynman theorem [1] directly connects the accuracy of the charge density to the forces acting on the nuclei, emphasizing its importance for determining molecular geometries and molecular dynamics. Additionally, the Kohn-Sham theorems [2, 3], which form the basis of density functional theory (DFT), prove that the ground state electron density uniquely determines ground state properties. Beyond these theoretical principles, charge densities are modeled for example in charge equilibration schemes used in polarizable force fields, where the dynamic redistribution of charges is the key to determining molecular interactions, polarization effects, and reactivity in complex systems.
@@
In recent years, machine learning potentials have become an important addition to ab initio methods, reproducing electronic structure calculations with high accuracy at only a fraction of the computational cost [4, 5, 6, 7, 8, 9, 10, 11]. Many recent machine learning potentials [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] include effects like electrostatics and long range charge transfer, which necessitates the ability to accurately model charge densities using ab initio methods. Notably, some approaches [23, 24, 25, 26] focus on directly modeling the electron density itself, using it for constructing transferable and interpretable potentials. These advancements highlight the critical role of accurate charge densities. Due to the high computational costs it is generally not feasible to compute reference data such as atomic Hirshfeld charges [27] using highly accurate methods like coupled cluster with single and double excitations (CCSD) [28, 29]. Instead, more approximate methods like DFT [2, 3] are typically used, trading some accuracy for increased computational efficiency. However, the accuracy of charge densities is strongly influenced by the choice of the exchange correlation functional and even the most accurate DFT functionals suffer from an electron self-interaction error [30, 31, 32, 33] that impacts the accuracy of the charge density. Medvedev et al. [34] even claim that after the early 2000s the density error in DFT functionals is in fact increasing. Additionally, several studies [35, 36, 37, 38] have analyzed atomic charge densities to identify and understand the error sources in DFT, focusing on the impact of functionals and electron self interaction errors. Controlling the self interaction error in approximate forms of the exchange correlation functional is a highly active field of research, with many approaches aiming to remove this source of error [39, 40, 41, 42, 43].
@@ II Methods
II.1 Hirshfeld charges
Assigning partial charges to atoms in molecules is an important task in computational chemistry, as these charges serve several purposes: they can be used as parameters in force fields and machine learned potentials, or to qualitatively analyze the electronic structure of the system under study. Hirshfeld charges are one of many methods [51, 52, 53, 54, 55, 56, 57] to partition the electron density into atomic partial charges. The Hirshfeld charge [27] of atom is defined as
Hirshfeld charges are typically used to analyze partial charges of atoms in molecules. However, there are some concerns that Hirshfeld charges are not well suited for this task since for example the dipole and quadrupole moments of the Hirshfeld charges differ significantly from the correct moments of the unpartitioned charge density [58], like those of many other partitioning methods. Hirshfeld charges can also be used to measure the dissimilarities of charge densities. In this work, one measure for the quality of a functional’s charge density is the deviation of its Hirshfeld charges from those obtained using a CCSD reference. Hirshfeld charges are also used in several third- and fourth-generation machine learning potentials [59, 12, 13, 14] where they provide the reference data for modeling long range electrostatics and charge transfer.
Ratios of two very small quantities can be evaluated numerically in a numerically stable way with the log-sum-exp-trick. The log-sum-exp (LSE) operation is defined as . By shifting all components wit , the LSE can be evaluated accurately even if the range of is large: . Let . Using the numerically stable LSE function, the partitioning function can be rewritten as
@@ II.3 Multi resolution wavelets
The adaptivity of multi wavelets allows to perform all-electron calculations where the resolution is increased for rapidly varying core orbitals, focusing computational resources where needed. Rigorous error control within multi wavelet methods enables the user to define the desired precision before the calculation begins, ensuring that the basis set error remains smaller than the chosen precision. For low precision wavelet calculations are generally slower than those using Gaussian orbitals, but when a large number of Gaussians is employed to obtain high precision, their efficiency becomes comparable. These features make multi wavelets particularly well suited for reference calculations in quantum mechanical applications, where both accuracy and efficiency are essential.
@@
The multi wavelet approach can be coupled by a separable form to represent Greens functions of the Poisson and Helmholtz equation [66] which enables efficient Hartree Fock [65] and DFT [64] calculations. These features are implemented in the software package MRChem [60, 61] and this code will be used for all wavelet based calculations in this paper.
III Results
III.1 Basis set errors of charge densities
which is a one dimensional function that contains the charge inside a sphere of radius . When spherically symmetric charge densities are considered, no information about the three dimensional charge density is lost. When comparing the charge density obtained with a DFT functional and a CCSD charge density, plotting the difference in offers insights into the spatial error distribution of the charge density. The difference in between two methods has units of charge and is the difference of charge contained in a sphere with radius . It is called when a DFT charge density is compared to CCSD results and when a DFT charge density that was obtained using a Gaussian basis set is compared with a multi wavelet calculation in this paper.
First, the error of the charge density for polarization consistent Gaussian basis sets [45] that are optimized for DFT calculations is analyzed. In Fig. 1, the difference of between an basis-set-error-free multi wavelet charge density and a charge density obtained with polarization consistent basis sets [45] is shown. The charge density was calculated with the PBE functional for a single neon atom. The “aug” prefix indicates the addition of diffuse functions in the Gaussian basis set. The aug-PC3 and aug-PC4 basis sets agree well with the multi wavelet reference calculation. For comparison, a charge density that was obtained using a correlation consistent basis set (aug-cc-pV5Z) is added in Fig. 1 which gives the same level of accuracy in the charge density. The aug-PC3 contains fewer basis functions than the aug-cc-pV5z basis set, the aug-PC4 more. In Fig. 2 the basis set error of charge densities obtained with a correlation consistent basis [46] sets is shown. Also the aug-cc-pV5Z basis set matches the multi wavelet result almost perfectly. It is therefore recommended to use this basis set whenever highly accurate charge densities are necessary with wave function based methods. The aug-PC3 and aug-PC4 offer the same level with a similar number of basis functions, the aug-PC3 is even a bit smaller than the aug-cc-pV5Z and is therefore an interesting alternative to the correlation consistent basis set when DFT is used. In Fig. 3 it is shown that the aug-cc-pV5Z basis set also produces highly accurate charge densities for coupled cluster calculations with single and double excitations. The error of charge densities with the cc-pVNZ and the PC basis is analyzed in Fig. 4 for a methanol molecule. The reference charge density was obtained using a multi wavelet DFT calculation with the PBE functional. As a measure of error, in this case the absolute value of the difference in Hirshfeld charges is used. Here, the addition of diffuse Gaussian basis functions greatly improves the accuracy and once again, the aug-cc-pV5Z basis is able to produce highly accurate charge densities. The aug-PC3 and aug-PC4 basis sets also produce highly converged Hirshfeld charges, the aug-PC4 appears to be the most accurate basis set in this comparison. However, the calculation using the aug-PC4 basis set was difficult to converge since the condition number of the overlap matrix from the basis functions was larger than . The aug-PC3, aug-PC4 and the aug-cc-pV5Z bases offer sufficient precision to accurately compare charge densities obtained with different functionals. Given that DFT results are being compared to CCSD calculations, it is more consistent to use the correlation-consistent aug-cc-pV5Z basis set, which is specifically designed for post-Hartree–Fock methods. Therefore, in the remainder of this paper, the aug-cc-pV5Z basis set will be used for comparing various exchange-correlation functionals.
III.2 Accuracy measures of charge densities
@@
Having found a suitable basis set to calculate high-precision charge densities, the charge densities for a variety of different exchange correlation functionals can be compared with accurate coupled cluster calculations with single and double excitations (CCSD). PySCF [44] was used for the CCSD calculations and the DFT calculation employing Gaussian basis sets. The charge densities of twelve different DFT functionals and Hartree Fock are benchmarked against CCSD charge densities for several molecules that are small enough to allow a CCSD calculation with an aug-cc-pV5Z basis set. The molecules are:
@@
Root mean squared error (RMSE) of the Hirshfeld charges
@@
Average infinity norm of difference in dipole moments
@@
Average infinity norm of difference in quadrupole moments
@@
The integrals required to calculate the norms were evaluated numerically with the integration grids of PySCF using the accurate grid level. The spherically symmetric reference charge densities required for the calculation of the Hirshfeld charges were computed using Abinit [68]. In Fig. 6, these measures are shown for all functionals. The functionals tested are:
III.3 Accuracy of different exchange correlation functionals
@@
We now analyse the accuracy of the different exchange correlation functionals for the investigated set of small molecules. Our previously defined measures of accuracy of the charge densities can be divided in two categories: point-wise measures and integral-based measures. The only point-wise measure used is the average maximal difference in charge density. The maximal difference in charge density is always located at the nuclear cusps. In this point-wise comparison, the Hartree Fock charge density most accurately approximates the CCSD charge density. The LDA charge density is worst regarding the point-wise measure and there is no clear trend in the remaining DFT functionals.
@@
However, the situation changes when considering the integral-based measures. In this case, HF and LDA yield poor charge densities, while all other functionals show significantly higher accuracy. The PBE, rPBE, and the BLYP functional are able to reproduce CCSD charge densities accurately in most tests but they struggle to accurately predict good dipole and quadrupole moments. The functionals that consistently produce good charge densities all either hybrid or meta-GGA. There is no clear benefit in the usage of range-separated hybrid functionals compared to modern state-of-the-art meta-GGA functionals such as r2-SCAN. The accuracy in the charge densities is similar in meta-GGA and hybrid functionals, making meta-GGA functionals the optimal choice that balance accuracy in charge densities and computational cost. Moreover, exact exchange is not needed to compute accurate Hirshfeld charges. The SCAN, r-SCAN and r2-SCAN functionals produce highly accurate Hirshfeld charges with an average relative error of around 4 percent making them ideal methods to compute reference data for machine learning potentials. B3LYP, PBE0, SCAN0 and the range-separated hybrid functionals CAM-B3LYP and WB97X-V are similarly accurate. Since they all contain exact exchange they are computationally significantly more expensive.
@@
Very similar trends are found when analysing a single atom. In Fig. 5, the difference in is shown for the tested functionals. It is found that r2-SCAN, SCAN and SCAN0 have the smallest error closely followed by PBE0, B3LYP and and the range-separated hybrid functional WB97X-V.
@@ IV Conclusion
A thorough investigation of the charge density error of Gaussian basis sets revealed that a very large basis set is required to accurately model the charge density. Polarization consistent basis sets that have been originally optimized for DFT calculations contain lower angular momenta in the Gaussian basis sets compared to the correlation consistent basis sets optimized for post-Hartree Fock calculations. The aug-cc-pV5Z, aug-PC3 and aug-PC4 basis sets are able to obtain accurate DFT charge densities making them the optimal choice for situations where highly accurate charge densities are needed. When molecular systems are considered, adding diffuse basis functions greatly improves the accuracy of the charge density. The aug-cc-pV5Z basis set appears to be accurate enough to obtain good charge densities. Our results are also of interest in the context of constructing machine learned potentials whose atomic charges are fitted to density functional results [12]. Since the atomic Hirshfeld charges obtained by different exchange correlation functionals differ by a few percent, there is no point trying to fit them in machine learning schemes to much higher precision.
An analysis of various metrics assessing the charge density quality showed that modern DFT functionals outperform Hartree-Fock for simple molecules. In these cases, density corrected DFT calculations [86] will therefore not improve the accuracy. The only functional that performs worse than HF is LDA. In contrast, all GGA, meta-GGA, and hybrid functionals evaluated in this study yield more accurate charge densities than HF. This trend suggests that the modern variants of meta-GGA and hybrid functionals are continuously enhancing the accuracy of charge densities in electronic structure calculations. The sole measure where HF excels is the maximal point-wise error in the charge density, which consistently occurs at the nuclear positions.

### v4

@@ Abstract
Accurate charge densities are essential for reliable electronic structure calculations because they significantly impact predictions of various chemical properties and in particular, according to the Hellmann-Feynman theorem, atomic forces. This study examines the accuracy of charge densities obtained from different DFT exchange-correlation functionals in comparison with coupled cluster calculations with single and double excitations. We find that modern DFT functionals can provide highly accurate charge densities, particularly in case of meta-GGA and hybrid functionals. In connection with Gaussian basis sets, it is necessary to use the largest basis sets available to obtain densities that are nearly basis set error free. These findings highlight the importance of selecting appropriate computational methods for generating high-precision charge densities, which are for instance needed to generate reference data for training modern machine learned potentials.
@@ I Introduction
Obtaining accurate charge densities is essential in any contetxt where quantum mechanical calculations are required. According to the Hohenberg-Kohn theorem [1], the charge density contains the same complete information as the much more complicated many-body wave function. In particular, it states that the ground state electron density uniquely determines the ground state and all its properties. The Hellmann-Feynman theorem [2] connects the charge density to the forces acting on the nuclei. Since integrating the scalar product between the forces and the atomic displacements gives energy differences, the charge density is in this way directly related to the energy differences. In addition forces are essential quantities in many types of simulations such as geometry relaxations and molecular dynamics.
Beyond these theoretical principles, charge densities are modeled for example in charge equilibration schemes used in polarizable force fields, where the dynamic redistribution of charges is the key to determining molecular interactions, polarization effects, and reactivity in complex systems.
@@
In recent years, machine learning potentials have become an important addition to ab initio methods, reproducing electronic structure calculations with high accuracy at only a fraction of the computational cost [3, 4, 5, 6, 7, 8, 9, 10]. Many recent machine learning potentials [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] include effects like electrostatics and long range charge transfer, which necessitates the ability to accurately model charge densities using ab initio methods. Notably, some approaches [22, 23, 24, 25] focus on directly modeling the electron density itself, using it for constructing transferable and interpretable potentials. These advancements highlight the critical role of accurate charge densities. Due to the high computational costs it is generally not feasible to compute reference data such as atomic Hirshfeld charges [26] using highly accurate methods like coupled cluster with single and double excitations (CCSD) [27, 28]. Instead, more approximate methods like DFT [1, 29] are typically used, trading some accuracy for increased computational efficiency. However, the accuracy of charge densities is strongly influenced by the choice of the exchange correlation functional and even the most accurate DFT functionals suffer from an electron self-interaction error [30, 31, 32, 33] that impacts the accuracy of the charge density. Medvedev et al. [34] analyzed electron densities for isolated atoms produced by 128 different DFT functionals and observed that while these densities became progressively closer to the exact ones until the early 2000s, this trend was reversed with the advent of unconstrained functionals that prioritized empirical fitting over physical rigor. Additionally, several studies [35, 36, 37, 38] have analyzed atomic charge densities to identify and understand the error sources in DFT, focusing on the impact of functionals and electron self interaction errors. Controlling the self interaction error in approximate forms of the exchange correlation functional is a highly active field of research, with many approaches aiming to remove this source of error [39, 40, 41, 42]. By evaluating semi-local density functionals on Hartree-Fock densities for transition states, Kaplan et al. [43] demonstrated that this approach improves the accuracy barrier heights obtained with DFT calculations. They attributed this improvement not to a reduction in density-driven errors, but to the introduction of a large density-driven error that cancels the functional-driven error which is consistent with our findings.
@@ II Methods
II.1 Kohn-Sham density functional theory
Kohn-Sham DFT can be categorized into different approximations for the exchange correlation functional. The most common used approximations are:
Local density approximation (LDA) where is a functional of the electron density at the same point.
Generalized gradient approximation (GGA) where is a functional of the electron density and its gradient.
meta GGA (mGGA) where is a functional of the electron density, its gradient and the kinetic energy density.
Hybrid functionals use a linear combination of an exchange correlation functional (either LDA, GGA or mGGA) and the Hartree-Fock exchange energy.
Range-separated hybrid functionals, in which the kernel in the exchange energy is screened by a parameter . The screening is applied through the expression , where denotes the error function.
@@ II.3 Multi resolution wavelets
The adaptivity of multi wavelets allows performing all-electron calculations where the resolution is increased for rapidly varying core orbitals, focusing computational resources where needed. Rigorous error control within multi wavelet methods enables the user to define the desired precision before the calculation begins, ensuring that the basis set error remains smaller than the chosen precision. For low precision wavelet calculations are generally slower than those using Gaussian orbitals, but when a large number of Gaussians is employed to obtain high precision, their efficiency becomes comparable. These features make multi wavelets particularly well suited for reference calculations in quantum mechanical applications, where both accuracy and efficiency are essential.
@@
The multi wavelet approach can be coupled by a separable form to represent Greens functions of the Poisson and Helmholtz equation [55] which enables efficient Hartree-Fock [54] and DFT [53] calculations. These features are implemented in the software package MRChem [56, 57] and this code will be used for all wavelet based calculations in this paper.
II.4 Accuracy measures of charge densities
@@
The charge densities for a variety of different exchange correlation functionals were compared with accurate coupled cluster calculations with single and double excitations (CCSD). PySCF [44] was used for the CCSD calculations and the DFT calculation employing Gaussian basis sets. The charge densities of twelve different DFT functionals and Hartree-Fock are benchmarked against CCSD charge densities for several molecules that are small enough to allow a CCSD calculation with an aug-cc-pV5Z basis set. In Section III.1, the precision of different Gaussian basis sets is compared and the choice of the aug-cc-V5Z basis set is justified. The molecules are:
@@
Average difference of the Hirshfeld charges
@@
Average component-wise difference in dipole moments
@@
Average component-wise difference in quadrupole moments
@@
The integrals required to calculate the norms were evaluated numerically with the integration grids of PySCF using the accurate grid level. The spherically symmetric reference charge densities required for the calculation of the Hirshfeld charges were computed using Abinit [58].
Measures 2–4 represent norms of the charge density error, where a value of 0 indicates a perfect match and higher values indicate larger errors. Measures 1, 5, and 6 are physically motivated and are best interpreted by comparing the results across the different functionals. To illustrate the significance of deviations in Hirshfeld charges, consider a hydrogen molecule with a bond length of 1.4 Bohr. A change in the Hirshfeld charge by would correspond to an electron-nucleus interaction energy shift equal to the chemical accuracy threshold of 1 kcal/mol. As a result, even relatively small changes in Hirshfeld charges can significantly impact the total energy.
In Fig. 6, these measures are shown for all functionals. The functionals tested are:
II.5 Hirshfeld charges
Assigning partial charges to atoms in molecules is an important task in computational chemistry, as these charges serve several purposes: they can be used as parameters in force fields and machine learned potentials, or to qualitatively analyze the electronic structure of the system under study. In this work, Hirshfeld charges are used as a measure for the quality of a functional’s charge density by calculating the deviation to Hirshfeld charges obtained using CCSD reference calculations. Hirshfeld charges are one of many methods [77, 78, 79, 80, 81, 82, 83] to partition the electron density into atomic partial charges. The Hirshfeld charge [26] of atom is defined as
Ratios of two very small quantities can be evaluated numerically in a numerically stable way with the log-sum-exp-trick. The log-sum-exp (LSE) operation is defined as . By shifting all components with , the LSE can be evaluated accurately even if the range of is large: . Let . Using the numerically stable LSE function, the partitioning function can be rewritten as
III Results
III.1 Basis set errors of charge densities
which is a one dimensional function that contains the charge inside a sphere of radius . When spherically symmetric charge densities are considered, no information about the three-dimensional charge density is lost. When comparing the charge density obtained with a DFT functional and a CCSD charge density, plotting the difference in offers insights into the spatial error distribution of the charge density. The difference in between two methods has units of charge and is the difference of charge contained in a sphere with radius . It is called when a DFT charge density is compared to CCSD results and when a DFT charge density that was obtained using a Gaussian basis set is compared with a multi wavelet calculation in this paper.
First, the error of the charge density for polarization consistent Gaussian basis sets [45] that are optimized for DFT calculations is analyzed. In Fig. 1, the difference of between a basis-set-error-free multi wavelet charge density and a charge density obtained with polarization consistent basis sets [45] is shown. The charge density was calculated with the PBE functional for a single neon atom. The “aug” prefix indicates the addition of diffuse functions in the Gaussian basis set. The aug-PC3 and aug-PC4 basis sets agree well with the multi wavelet reference calculation. For comparison, a charge density that was obtained using a correlation consistent basis set (aug-cc-pV5Z) is added in Fig. 1 which gives the same level of accuracy in the charge density. The aug-PC3 contains fewer basis functions than the aug-cc-pV5z basis set, the aug-PC4 more. In Fig. 2 the basis set error of charge densities obtained with a correlation consistent basis [46] sets is shown. Also, the aug-cc-pV5Z basis set matches the multi wavelet result almost perfectly. It is therefore recommended using this basis set whenever highly accurate charge densities are necessary with wave function based methods. The aug-PC3 and aug-PC4 offer the same level with a similar number of basis functions, the aug-PC3 is even a bit smaller than the aug-cc-pV5Z and is therefore an interesting alternative to the correlation consistent basis set when DFT is used. In Fig. 3 it is shown that the aug-cc-pV5Z basis set also produces highly accurate charge densities for coupled cluster calculations with single and double excitations.
To investigate the basis set error of charge densities for a system without spherical symmetry, Hirshfeld charges are analyzed in Fig. 4 for a methanol molecule. The reference charge density was obtained using a multi wavelet DFT calculation with the PBE functional. As a measure of error, in this case the absolute value of the difference in Hirshfeld charges is used. Here, the addition of diffuse Gaussian basis functions greatly improves the accuracy and once again, the aug-cc-pV5Z basis is able to produce highly accurate charge densities. The aug-PC3 and aug-PC4 basis sets also produce highly converged Hirshfeld charges, the aug-PC4 appears to be the most accurate basis set in this comparison. However, the calculation using the aug-PC4 basis set was difficult to converge since the condition number of the overlap matrix from the basis functions was larger than . The aug-PC3, aug-PC4 and the aug-cc-pV5Z bases offer sufficient precision to accurately compare charge densities obtained with different functionals.
Given that DFT results are being compared to CCSD calculations, it is more consistent to use the correlation-consistent aug-cc-pV5Z basis set, which is specifically designed for post-Hartree-Fock methods.
In Fig. 4 of our recent work on accurate force calculations using multi-wavelets [57], we analyzed the energy and force convergence for the same methanol molecule used in this study to compute Hirshfeld charges. That analysis showed that achieving chemical accuracy requires at least a quadruple-zeta basis set. To ensure an additional margin of accuracy, we opted for a quintuple-zeta basis set. Moreover, the earlier study demonstrated that augmented basis functions significantly improve the accuracy of forces. A trend similarly observed for the Hirshfeld charges in this work. Therefore, in the remainder of this paper, the aug-cc-pV5Z basis set will be used for comparing various exchange-correlation functionals.
III.2 Accuracy of different exchange correlation functionals
@@
We now analyze the accuracy of the different exchange correlation functionals for the investigated set of small molecules. Our previously defined measures of accuracy of the charge densities can be divided in two categories: point-wise measures and integral-based measures. The only point-wise measure used is the average maximal difference in charge density. The maximal difference in charge density is always located at the nuclear cusps. In this point-wise comparison, the Hartree-Fock charge density most accurately approximates the CCSD charge density. The LDA charge density is worst regarding the point-wise measure and there is no clear trend in the remaining DFT functionals.
@@
However, the situation changes when considering the integral-based measures. In this case, HF and LDA yield poor charge densities, while all other functionals show significantly higher accuracy. The PBE, rPBE, and the BLYP functional are able to reproduce CCSD charge densities accurately in most tests, but they struggle to accurately predict good dipole and quadrupole moments. The functionals that consistently produce good charge densities all either hybrid or meta-GGA. There is no clear benefit in the usage of range-separated hybrid functionals compared to modern state-of-the-art meta-GGA functionals such as r2-SCAN. The accuracy in the charge densities is similar in meta-GGA and hybrid functionals, making meta-GGA functionals the optimal choice that balance accuracy in charge densities and computational cost. Moreover, exact exchange is not needed to compute accurate Hirshfeld charges. The SCAN, r-SCAN and r2-SCAN functionals produce highly accurate Hirshfeld charges with an average relative error of around 4 percent making them ideal methods to compute reference data for machine learning potentials. B3LYP, PBE0, SCAN0 and the range-separated hybrid functionals CAM-B3LYP and WB97X-V are similarly accurate. Since they all contain exact exchange they are computationally significantly more expensive. It is noteworthy that all empirically designed functionals exhibit a significantly larger spread in at least one error measure compared to the other methods. In contrast, the strongly constrained SCAN and r2-SCAN functionals, which satisfy all known meta-GGA functional constraints, consistently show less variation. The r-SCAN functional exhibits a significantly larger spread in the average quadrupole difference compared to the other functionals, possibly due to its violation of some known meta-GGA constraints [72].
@@
Very similar trends are found when analyzing a single atom. In Fig. 5, the difference in is shown for the tested functionals. It is found that r2-SCAN, SCAN and SCAN0 have the smallest error closely followed by PBE0, B3LYP and the range-separated hybrid functional WB97X-V.
@@ IV Conclusion
A thorough investigation of the charge density error of Gaussian basis sets revealed that a very large basis set is required to accurately model the charge density. Polarization consistent basis sets that have been originally optimized for DFT calculations contain lower angular momenta in the Gaussian basis sets compared to the correlation consistent basis sets optimized for post-Hartree-Fock calculations. The aug-cc-pV5Z, aug-PC3 and aug-PC4 basis sets are able to obtain accurate DFT charge densities making them the optimal choice for situations where highly accurate charge densities are needed. When molecular systems are considered, adding diffuse basis functions greatly improves the accuracy of the charge density. The aug-cc-pV5Z basis set appears to be accurate enough to obtain good charge densities. Our results are also of interest in the context of constructing machine learned potentials whose atomic charges are fitted to density functional results [11]. Since the atomic Hirshfeld charges obtained by different exchange correlation functionals differ by a few percent, there is no point trying to fit them in machine learning schemes to much higher precision.
An analysis of various metrics assessing the charge density quality showed that modern DFT functionals outperform Hartree-Fock for simple, mostly sp-bonded molecules at their equilibrium geometries. For systems containing stretched bonds, or orbitals with higher angular momenta such as d and f orbitals, further benchmarking will be required. Conclusions from this analysis are therefore valid for similar systems. The only functional that performs worse than HF is LDA. In contrast, all GGA, meta-GGA, and hybrid functionals evaluated in this study yield more accurate charge densities than HF. This trend suggests that the modern variants of meta-GGA and hybrid functionals are continuously enhancing the accuracy of charge densities in electronic structure calculations. The sole measure where HF excels is the maximal point-wise error in the charge density, which consistently occurs at the nuclear positions.
Notably, the functionals that adhere to theoretical constraints, such as SCAN and r2-SCAN, produced the most consistent and accurate results across all error measures. In contrast, empirically designed functionals exhibited a significantly larger spread in at least one error measure, emphasizing the importance of adhering to physically rigorous constraints for achieving consistently accurate charge densities.